Karlottalotta:

As shown below, an observer (O) is located 660 feet from a tree (T). The observer notices a hawk (H) flying at a 35° angle of elevation from his line of sight. What equation and trigonometric function can be used to solve for the height (h) of the hawk? What is the height of the hawk? You must show all work and calculations to receive full credit

Karlottalotta:

1 attachment
snowflake0531:

\[sin(\theta)=\frac{opposite}{hypotenuse}\\cos(\theta)=\frac{adjacent}{hypotenuse}\\tan(\theta)=\frac{opposite}{adjacent}\] Looking from the reference angle, the 35 degrees one We want to find opposite and we have adjacent Which trig ratio should we use?

surjithayer:

\[\tan 35=\frac{ h }{ 660 }\] \[h=660\tan 35=?\]